OpenStudy (anonymous):

Evaluate the integral ∫_0^1{36dx/((2x+1)^3)}

OpenStudy (anonymous):

\[\int\limits_{0}^{1}{\frac{36dx}{(2x+1)^3}}\]

OpenStudy (anonymous):

Okay: \[\int\limits_0^1 \frac{36dx}{(2x+1)^3}\] The first thing you want to do is make a usubstitution. Let u=2x+1 then du=2dx or dx=(1/2)du Re-evaluate your limits plugging in zero you get 1 and plugging in 1 your get 3. So you have: \[\frac{36}{2}\int\limits_1^3\frac{du}{u^3}\] Rewrite it as: \[18 \int\limits_1^3 u^{-3}du=18[\frac{-1}{2}u^{-2}]_1^3=-9(\frac{1}{3^2}-\frac{1}{1^2})=-9*\frac{-8}{9}=8\]

OpenStudy (anonymous):

ok so i know my u=2x+1 du=2 18du=36 18∫u^3du 18u^4/4 (18(2x+1)^4/4) 18(2(1)+1)^4/4-18(2(0)+1)^4/4 364.5-4.5

OpenStudy (anonymous):

It becomes u^-3. Not u^3 :P

OpenStudy (anonymous):

ok i see where i went wrong thank yu malevolence

OpenStudy (anonymous):

No problem :)